# Estimating the FDR of variable selection

### v2

@@ 4.1 HIV drug resistance studies
Our initial processing of the data follows that of Barber and Candès (2016), yielding a different regression data set for each of 16 drugs in three different drug classes. In the data set for a given drug, the response $y_{i}$ is a measure of sample $i$’s resistance to that drug, and the explanatory variables are binary indicators of whether each of several hundred mutations was present or absent in that sample.
@@
Figure 2 shows our FDR estimates alongside the CV estimate of MSE, for four selected experiments representing a variety of statistical scenarios. Results for all drugs are available in Appendix E. The error bars for both CV and $\widehat{\textnormal{FDR}}$ our estimator show $\pm$±one standard error of $\widehat{\textnormal{FDR}}$ estimated by bootstrap. The $\widehat{\textnormal{FDP}}$ curves in the figures are estimated using the TSM signals as an approximation to the ground truth.
The four panels of Figure 2 illustrate how different variable selection problems can have very different trade-offs between variable selection accuracy and prediction accuracy. These scenarios would be impossible to distinguish from inspecting the CV curves, but our $\widehat{\textnormal{FDR}}$ estimator is quite effective in distinguishing them from each other. Panel (a) shows an experiment in which variable selection appears to be easy: the one-standard-error rule simultaneously achieves good model fit and low FDR. Panel (b), by contrast, shows a regime in which no lasso estimator can simultaneously achieve competitive model fit and low FDR. Panel (c) shows an experiment in which the estimated FDR is non-monotone in the complexity parameter $\lambda$: both our estimator and the TSM $\widehat{\textnormal{FDP}}$ judge the fifth variable to enter the model, gene P71.V, to be a likely noise variable, because its $t$-statistic is $-0.38$, corresponding to a $p$-value of $0.70$, and it is also absent from the TSM list (the $t$-statistics for the first four variables to enter all exceed 5). Panel (d) shows one of the three experiments in which our $\widehat{\textnormal{FDR}}$ departs substantially from the TSM $\widehat{\textnormal{FDP}}$. In this experiment, our $\widehat{\textnormal{FDR}}$ estimator is large mainly because three of the first five selected variables have $p$-values equal to $0.77$, $0.53$, and $0.18$, but the TSM $\widehat{\textnormal{FDP}}$ curve is small because all three variables appear in the TSM list.
@@ 4.2 Protein-signaling network
Figure 4 shows two representative examples of the results from running the graphical Lasso on the $14$ experimental settings. Results for all experiments are available in Appendix E. As in the HIV study, we show our $\widehat{\textnormal{FDR}}$ and its estimated standard error, alongside an estimate of FDP calculated using the causal DAG supplied by Sachs et al. (2005) (denoted as the DAG $\widehat{\textnormal{FDP}}$ in Figure 4), and the CV estimate of the predictive log-likelihood. In both panels of Figure 4, the models that we would select by cross-validation are nearly useless: one selects almost no edges, while the other selects every edge. Friedman et al. (2008) analyzed the data somewhat differently than we did, pooling all $14$ experiments together, but likewise reported that the CV curve continuously decreased until all edges were present in the model. By contrast, our $\widehat{\textnormal{FDR}}$ gives useful information about the selection accuracy, and suggests in each experiment a natural stopping point.
@@ 5.1 Gaussian linear model
[…]
@@
[…]
@@
[…]
@@
Figure 5: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ in the Gaussian linear model. The error bars show the $5\%$ to $95\%$ quantiles of the empirical distributions and the solid lines show the sample average. $\widehat{\textnormal{FDR}}$ shows a small non-negative bias in estimating FDR and $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$ successively captures the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$.
@@ 5.2 Model-X setting
Figure 7: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ in the model-X settings. The error bars show the $5\%$ to $95\%$ quantiles of the empirical distributions and the solid lines show the sample average. $\widehat{\textnormal{FDR}}$ shows a small non-negative bias in estimating FDR and $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$ successively captures the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$.
@@ 5.3 Gaussian graphical model
Figure 8: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ in the Gaussian graphical model. The error bars show the $5\%$ to $95\%$ quantiles of the empirical distributions and the solid lines show the sample average. $\widehat{\textnormal{FDR}}$ shows a small non-negative bias in estimating FDR and $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$ successively captures the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$.
@@ 5.4 Unfavorable case: equicorrelated Gaussian
Figure 9: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ on the MCC problem. The error bars show the $5\%$ to $95\%$ quantiles of the empirical distributions and the solid lines show the sample average. $\widehat{\textnormal{FDR}}$ shows a large variance but $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$ successively captures the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$.
@@ 5.5 Robustness to model misspecification
[…]
@@ 6 Computational details
for $j=1,\ldots,d$ with $\psi_{j}\neq 0$. These calculations can be easily parallelized. In a generic setting, we can use Monte-Carlo integration to approximate $\widehat{\textnormal{FDR}}^{*}_{j}$. The worst-case time complexity for computing $\widehat{\textnormal{FDR}}$ is $O(d\cdot M\cdot T)$, where $M$ is the number of Monte-Carlo samples and $T$ is the running time of the selection procedure. We note that calculating the integrand of $\widehat{\textnormal{FDR}}_{j}^{*}$ for each Monte-Carlo sample has the same computation complexity as calculating one placebo statistic in the conditional randomization test (CRT, Candès et al. (2018)) with the same selection procedure for variable $j$. In our implementation, we use $20$ Monte-Carlo samples for computing each $\widehat{\textnormal{FDR}}^{*}_{j}$ and $10$ Bootstrap samples.
@@
The running time of evaluating $\widehat{\textnormal{FDR}}$ (without bootstrap s.e. estimation) at all $10$ different tuning parameter values in the simulation problems in Section 5 is typically several minutes on a single-cored computer. We report the exact running time in Appendix F.
7 Summary
@@
Our FDR estimator is implemented in an R package available online at the Github repository
github.com/yixiangLuo/hFDR. The R code to reproduce all simulations and figures is available at github.com/yixiangLuo/hFDR_expr.
@@ A.1 Simulation settings for the illustrative examples
[…]
@@
In this section, we describe the cross-validation algorithm we use to construct the set $\widehat{\mathcal{H}_{0}}$ in computing $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$. Formally, let $\boldsymbol{D}_{\text{trn},k}$ and $\boldsymbol{D}_{\text{val},k}$ be the $k$th training and validation data set in a $K$-folds cross-validation, respectively. Throughout the section, we write $\boldsymbol{\hat{\theta}}(\widehat{\mathcal{H}_{0}}(\lambda);\boldsymbol{D}_{\text{trn},k})$ with $\widehat{\mathcal{H}_{0}}(\lambda)=\mathcal{R}^{\mathsf{c}}(\lambda;\boldsymbol{D}_{\text{trn},k})$ as $\boldsymbol{\hat{\theta}}(\lambda;\boldsymbol{D}_{\text{trn},k})$ for simplicity of notation. We have
$\boldsymbol{\hat{\theta}}(\lambda;\boldsymbol{D}_{\text{trn},k})=\underset{\theta_{j}=0,\,j\not\in\mathcal{R}(\lambda;\boldsymbol{D}_{\text{trn},k})}{\operatorname*{argmax}}\;P_{\boldsymbol{\theta}}(\boldsymbol{D}_{\text{trn},k}).$
@@
[…]
@@
A.3 Sparse model estimation with graphical Lasso
@@ Appendix B Example where $\widehat{\textnormal{FDR}}$ has non-vanishing variance
=$\mathcal{R}=\left\{j:\;Z_{j}\geq 0.5\cdot\sqrt{2\log d}\right\},$
and set $\psi_{j}=\mathbf{1}\left\{p_{j}>0.5\right\}$, where $p_{j}$ is the one-sided marginal z-test p-value. Let $\boldsymbol{S}_{j}=\boldsymbol{Z}_{-j}-\boldsymbol{\Sigma}_{-j,j}\cdot Z_{j}$ and $\boldsymbol{\theta}=\boldsymbol{0}$. Figure 12 shows the standard deviation of $\widehat{\textnormal{FDR}}$ does not converge to 00 as the number of variables $d$ approach infinity.
Figure 12: For variable selection on strongly positively equicorrelated $Z$-variables, standard deviation of $\widehat{\textnormal{FDR}}$ does not converge to 00 as the number of variables approach infinity.
@@ Appendix C Fast algorithms for computing $\widehat{\textnormal{FDR}}^{*}_{j}$
$-\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{Y}-\boldsymbol{X}\cdot\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}})+\boldsymbol{\nu}=\boldsymbol{0},$
@@
Appendix D Deferred proofs
In this section, we assume Assumption 3.1 and prove the propositions in Section 3.1. Throughout the section, we use $\mathcal{R}(\lambda)$ to denote the selection set of Lasso. Define $\mathcal{R}_{k}=\mathcal{R}\cap B_{k}$ as the selected variables in block $k$ and $R_{k}=|\mathcal{R}_{k}|$ as its cardinality. Throughout the section, we consider any pair of $(j,k)$ such that $j\in B_{k}$. We start with showing two technical lemmas.

### v4

@@ 4.1 HIV drug resistance studies
Our initial processing of the data follows that of Barber and Candès (2016), yielding a different regression data set for each of 16 drugs in three different drug classes. In the data set for a given drug, the response $y_{i}$ is a measure of sample $i$’s resistance to that drug, and the explanatory variables are binary indicators of whether each of several hundred mutations was present or absent in that sample. As in Barber and Candès (2016), we adopt the Gaussian linear model assumption in this study.
@@
[…]
The four panels of Figure 2 illustrate how different variable selection problems can have very different trade-offs between variable selection accuracy and prediction accuracy. These scenarios would be impossible to distinguish from inspecting the CV curves, but our $\widehat{\textnormal{FDR}}$ estimator is quite effective in distinguishing them from each other. Panel (a) shows an experiment in which variable selection appears to be easy: the one-standard-error rule simultaneously achieves good model fit and low FDR. Panel (b), by contrast, shows a regime in which no lasso estimator can simultaneously achieve competitive model fit and low FDR. Panel (c) shows an experiment in which the estimated FDR is non-monotone in the complexity parameter $\lambda$: both our estimator and the TSM $\widehat{\textnormal{FDP}}$ judge the fifth variable to enter the model, gene P71.V, to be a likely noise variable, because its $t$-statistic is 0.38-0.38, corresponding to a $p$-value of $0.70$, and it is also absent from the TSM list (the $t$-statistics for the first four variables to enter all exceed 5). Panel (d) shows one of the three experiments in which our $\widehat{\textnormal{FDR}}$ departs substantially from the TSM $\widehat{\textnormal{FDP}}$. In this experiment, our $\widehat{\textnormal{FDR}}$ estimator is large mainly because three of the first five selected variables have $p$-values equal to $0.77$, $0.53$, and $0.18$, but the TSM $\widehat{\textnormal{FDP}}$ curve is small because all three variables appear in the TSM list.
@@ 4.2 Protein-signaling network
Figure 4 shows two representative examples of the results from running the graphical Lasso on the $14$ experimental settings. Results for all experiments are available in Appendix F. As in the HIV study, we show our $\widehat{\textnormal{FDR}}$ and its estimated standard error, alongside an estimate of FDP calculated using the causal DAG supplied by Sachs et al. (2005) (denoted as the DAG $\widehat{\textnormal{FDP}}$ in Figure 4), and the CV estimate of the predictive log-likelihood. In both panels of Figure 4, the models that we would select by cross-validation are nearly useless: one selects almost no edges, while the other selects every edge. Friedman et al. (2008) analyzed the data somewhat differently than we did, pooling all $14$ experiments together, but likewise reported that the CV curve continuously decreased until all edges were present in the model. By contrast, our $\widehat{\textnormal{FDR}}$ gives useful information about the selection accuracy, and suggests in each experiment a natural stopping point.
@@ 5.1 Gaussian linear model
    IID normal: independent Gaussian random variables $X_{j}\overset{\textnormal{i.i.d.}}{\sim}\mathcal{N}(0,1)$.
@@
    Sparse: Independent sparse binary observations $X_{ij}\overset{\textnormal{i.i.d.}}{\sim}\text{Bernoulli}(0.05)$.
@@
$\theta_{j}\overset{\textnormal{i.i.d.}}{\sim}\theta^{*}\cdot(1+\textnormal{Exp}(1))/2,$
@@
[…]
@@ 5.2 Model-X setting
[…]
@@ 5.3 Gaussian graphical model
[…]
@@ 5.4 Unfavorable case: equicorrelated Gaussian
[…]
@@ 5.5 Robustness to model misspecification
In the “t-noise” setting, we have the noise $\varepsilon$ in the Gaussian linear model follow the t-distribution with a degree of freedom $5$ instead of the standard normal distribution. The “Exponential” setting has the variables follow the heavy-tailed exponential distribution $X_{j}\overset{\textnormal{i.i.d.}}{\sim}\textnormal{Exp(1)}$ but are mistaken as Gaussian in the model-X assumption. Finally, In the “X Perturbation” setting, we multiply each observed variable by an IID uniformly distributed perturbation $U([0,2])$. As the figure shows, our $\widehat{\textnormal{FDR}}$ performs well in these scenarios.
@@ 6 Computational details
for $j=1,\ldots,d$ with $\psi_{j}\neq 0$. These calculations can be easily parallelized. In a general setting, $\widehat{\textnormal{FDR}}^{*}_{j}$ can be approximated via Monte Carlo integration: generate $M$ samples of the data $\boldsymbol{D}$ from the known conditional distribution, compute the integrand ${\mathbf{1}\left\{j\in\mathcal{R}\right\}}/{R}$ for each sample, and take the average over them. The worst-case time complexity for computing $\widehat{\textnormal{FDR}}$ is $O(d\cdot M\cdot T)$, where $T$ is the running time of the selection procedure. We note that calculating the integrand of $\widehat{\textnormal{FDR}}_{j}^{*}$ for each Monte-Carlo sample has the same computation complexity as calculating one placebo statistic in the conditional randomization test (CRT, Candès et al. (2018)) with the same selection procedure for variable $j$. In our implementation, we use $20$ Monte-Carlo samples for computing each $\widehat{\textnormal{FDR}}^{*}_{j}$ and $10$ Bootstrap samples.
@@
The running time of evaluating $\widehat{\textnormal{FDR}}$ (without bootstrap s.e. estimation) at all $10$ different tuning parameter values in the simulation problems in Section 5 is typically several minutes on a single-cored computer. We report the exact running time in Appendix G.
7 Summary and discussion
@@
While this paper is focused on three settings, our method can be applied insofar as a sufficient statistic $\boldsymbol{S}_{j}$ exists under the null submodel $H_{j}$, i.e., the data $\boldsymbol{D}$ has a known distribution conditional on $\boldsymbol{S}_{j}$ under $H_{j}$. Consider a common setting in the multiple testing literature where a test statistic $T_{j}$ is available for $H_{j}$ and $T_{1},\ldots,T_{d}$ are independent. If $H_{j}$ are all simple hypotheses (e.g., $H_{j}:\theta_{j}=0$), we can simply set $\boldsymbol{S}_{j}=(T_{1},\ldots,T_{j-1},T_{j+1},\ldots,T_{d}\}$. When the $H_{j}$ are composite but simple conditional on some auxiliary statistic $V_{j}$, we can set $\boldsymbol{S}_{j}=(T_{1},\ldots,T_{j-1},V_{j},T_{j+1},\ldots,T_{d}\}$. For example, if the researcher has two samples $X_{j1},\ldots,X_{jn_{j_{1}}}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}P$ and $Z_{j1},\ldots,Z_{jn_{j_{2}}}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Q$ for the hypothesis $H_{j}:P=Q$, $\boldsymbol{S}_{j}$ can be always chosen as the unordered set of $\{X_{j1},\ldots,X_{jn_{j_{1}}},Z_{j1},\ldots,Z_{jn_{j_{2}}}\}$. If $P=F_{\theta_{1}}$ and $Q=F_{\theta_{2}}$ for some exponential family $F$ with natural parameter $\theta$ and sufficient statistic $T(\cdot)$, then we can choose $\boldsymbol{S}_{j}=\sum_{i=1}^{j_{1}}T(X_{ji})+\sum_{k=1}^{j_{2}}T(Z_{jk})$. Therefore, our method can also be applied to a large class of multiple testing problems with independent test statistics.
@@
Our FDR estimator is implemented in an R package available online at the Github repository github.com/yixiangLuo/hFDR. The R code to reproduce all simulations and figures is available at github.com/yixiangLuo/hFDR_expr.
@@ A.1 Simulation settings for the illustrative examples
In Figure 1(a), the explanatory variables are independent Gaussian – $X_{j}\overset{\textnormal{i.i.d.}}{\sim}\mathcal{N}(0,1)$. We have $d=200$ variables, $n=600$ observations, and $d_{1}=20$ variables are signal variables with random signal strength $\theta_{j}=0.25\cdot(1+\textnormal{Exp}_{j}(1))/2$, where $\textnormal{Exp}_{j}(1)$ with $j\in\mathcal{H}_{0}^{\mathsf{c}}$ are independent exponentially distributed random variable with rate $1$. The response $Y$ is generated by the Gaussian linear model with $\sigma=1$.
@@
[…]
$\boldsymbol{\hat{\theta}}^{(*)}(\lambda;\boldsymbol{D}_{\text{trn},k})=\underset{\theta_{j}=0,\,j\not\in\mathcal{R}(\lambda;\boldsymbol{D}_{\text{trn},k})}{\operatorname*{argmax}}\;P_{\boldsymbol{\theta}}(\boldsymbol{D}_{\text{trn},k}).$
@@
$\lambda_{\text{cv}}=\underset{\lambda}{\operatorname*{argmax}}\;\frac{1}{K}\sum_{k=1}^{K}P_{\boldsymbol{\hat{\theta}}^{(*)}(\lambda;\boldsymbol{D}_{\text{trn},k})}(\boldsymbol{D}_{\text{val},k})$
@@
A.3 Bias contribution from signal variables
Theorem A.1.
In the Gaussian linear model (1), consider the canonical version (11) of $\widehat{\textnormal{FDR}}_{j}$:
$\widehat{\textnormal{FDR}}_{j}=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\cdot\frac{\mathbf{1}\{p_{j}>\zeta\}}{1-\zeta},$
with $p_{j}$ being the two-sided t-test p-value for $H_{j}$. Define $T_{j}$ as the t-test statistic with noncentrality
$\mu_{j}=\frac{\theta_{j}}{\sigma\sqrt{(\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1})_{jj}}}.$
Let $t_{1-\zeta/2}$ denote the $1-\zeta/2$ upper quantile of the central t-distribution with degree of freedom ndn-d. Then the contribution to the bias of $\widehat{\textnormal{FDR}}$ from variable $j$, when the signal strength $|\mu_{j}|>t_{1-\zeta/2}$, has upper bound
$\mathbb{E}[\widehat{\textnormal{FDR}}_{j}]\leq\frac{1}{(1-\zeta)}\,\left[2\exp\!\left(-\frac{\bigl(|\mu_{j}|-\sqrt{2}\,t_{1-\zeta/2}\bigr)_{+}^{2}}{2}\right)\;+\;\exp\!\left(-\frac{n-d}{16}\right)\right].$
Proof.
Note
$T_{j}\;\stackrel{{\scriptstyle d}}{{=}}\;\frac{Z+\mu_{j}}{\sqrt{U/(n-d)}},\qquad\textnormal{where }Z\sim\mathcal{N}(0,1),\ U\sim\chi^{2}_{n-d},\ Z\perp U.$
Let $A:=\{U\leq 2(n-d)\}$. Split
$\mathbb{P}(|T|<t_{1-\zeta/2})\leq\mathbb{P}(|T|<t_{1-\zeta/2},\ A)+\mathbb{P}(A^{c}).$
On $A$ we have $\sqrt{U/(n-d)}\leq\sqrt{2}$, hence
$\{|T|<t_{1-\zeta/2}\}\cap A\subseteq\{|Z+\mu_{j}|<\sqrt{2}\,t_{1-\zeta/2}\}.$
Therefore
$\mathbb{P}(|T|<t_{1-\zeta/2})\leq\mathbb{P}(|Z+\mu_{j}|<\sqrt{2}\,t_{1-\zeta/2})+\mathbb{P}(U>2(n-d)).$ (15)
For the Gaussian term, with $a:=\sqrt{2}\,t_{1-\zeta/2}$,
$\mathbb{P}(|Z+\mu_{j}|<a)=\mathbb{P}(-a-\mu_{j}<Z<a-\mu_{j})\leq\mathbb{P}(Z<a-\mu_{j})\leq\exp\!\left(-\frac{(\mu_{j}-a)_{+}^{2}}{2}\right),$
and applying the same bound to $\mu_{j}$ gives
$\mathbb{P}(|Z+\mu_{j}|<a)\leq 2\exp\!\left(-\frac{(|\mu_{j}|-a)_{+}^{2}}{2}\right).$
For the chi-square term, use a Chernoff bound,
$\mathbb{P}(U>2(n-d))\leq\exp\!\left(-\frac{n-d}{16}\right).$
Therefore
$\mathbb{P}(p_{j}>\zeta)=\mathbb{P}(|T_{j}|<t_{1-\zeta/2})\;\leq\;2\exp\!\left(-\frac{\bigl(|\mu_{j}|-\sqrt{2}\,t_{1-\zeta/2}\bigr)_{+}^{2}}{2}\right)\;+\;\exp\!\left(-\frac{n-d}{16}\right).$
Note
$\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\leq 1\qquad a.s.$
Combining all the components together, the inequality is proved. ∎
Theorem A.1 states that the contribution to the bias of $\widehat{\textnormal{FDR}}$ from strong signal variables decays to 0 exponentially with the signal strength. In the proof, the bound
$\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\leq 1\qquad a.s.$
is rather loose. In practice, we expect $R$ to be not much smaller than the number of strong signals for a reasonable variable selection procedure. Hence the bound established in Theorem A.1 is also a practical estimate for the bias coming from all the strong signals.
A.4 Sparse model estimation with graphical Lasso
@@ Appendix B Example where $\widehat{\textnormal{FDR}}$ has non-vanishing variance
$\mathcal{R}=\left\{j:\;Z_{j}\geq 0.5\cdot\sqrt{2\log d}\right\},$
and set $\psi_{j}=\mathbf{1}\left\{p_{j}>0.5\right\}$, where $p_{j}$ is the one-sided marginal z-test p-value. Let $\boldsymbol{S}_{j}=\boldsymbol{Z}_{-j}-\boldsymbol{\Sigma}_{-j,j}\cdot Z_{j}$ and $\boldsymbol{\theta}=\boldsymbol{0}$. Figure 12 shows the standard deviation of $\widehat{\textnormal{FDR}}$ does not converge to 0 as the number of variables $d$ approach infinity.
Figure 12: For variable selection on strongly positively equicorrelated $Z$-variables, standard deviation of $\widehat{\textnormal{FDR}}$ does not converge to 0 as the number of variables approach infinity.
@@ Appendix C Fast algorithms for computing $\widehat{\textnormal{FDR}}^{*}_{j}$
$\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{Y}-\boldsymbol{X}\cdot\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}})+\boldsymbol{\nu}=\boldsymbol{0},$
@@
Appendix D Variance of FDR estimation
In this section, we assume Assumption 3.1 and prove the propositions in Section 3.1. Throughout the section, we use $\mathcal{R}(\lambda)$ to denote the selection set of Lasso. Define $\mathcal{R}_{b}=\mathcal{R}\cap(b)$ as the selected variables in block $(b)$ and $R_{b}=|\mathcal{R}_{b}|$ as its cardinality. Throughout the section, we consider any pair of $(j,b)$ such that $j\in(b)$. We start with showing two technical lemmas.